# The Manakov-Zakharov-Ward model as an integrable decoupling limit of the membrane

### v2

@@ 1.1 The bosonic membrane
The action of the bosonic membrane (or equivalently the bosonic part of the supermembrane [2]) is
$S$ $\sim\int\mathrm{d}^{3}\sigma\left(\frac{1}{2}\sqrt{-\gamma}\gamma^{MN}\partial_{M}\hat{X}^{\mu}\partial_{N}\hat{X}^{\nu}g_{\mu\nu}\right$. (1.1)
in a bosonic supergravity background described by a metric $g$ and three-form gauge field $C$. $\sigma^{M}$ are the coordinates on the membrane world-volume, $\hat{X}^{\mu}(\sigma)$ the embedding coordinate fields, and $\gamma_{MN}$ is the (dynamical) world-volume metric. The cosmological constant term, as usual, ensures that the world-volume metric becomes the induced metric (as algebraic equation of motion for the world-volume metric $\gamma$):
$\gamma_{MN}=\partial_{M}\hat{X}^{\mu}\partial_{N}\hat{X}^{\nu}g_{\mu\nu}.$ (1.2)
${}\quad+\epsilon^{NKL}\frac{1}{3!}{F^{\mu}}_{\nu\kappa\lambda}\partial_{N}\hat{X}^{\nu}\partial_{K}\hat{X}^{\kappa}\partial_{L}\hat{X}^{\lambda},$ (1.3)
where $\Gamma$ denote Christoffel symbols and $F=\mathrm{d}C$. Due to the non-trivial nature of the world-volume metric $\gamma$, these equations are quite intractable, in general.
Famously [17, 18], even this supposedly most simple case leads to complications, such as a continuous spectrum and spike instabilities. Quantum non-integrability of the membrane in the sense of scattering of fluctuations around a classical solution, and its relation to string theory, has recently been shown in [19]. In this letter, we will see that a large wrapping limit together with a scaling limit of the background can simplify these equations. In the special case considered here, the result will even be a classically integrable system.
@@ 1.2 The Manakov-Zakharov-Ward model
The model is of $\sigma$-model type and can be formulated in terms of maps $g:\Sigma\rightarrow G$ from a $(1+2)$-dimensional flat spacetime $\Sigma$ (referred to as the world-volume) into a Lie group $G$. The dynamics is conveniently expressed using the left-invariant Maurer–Cartan current $j=g^{-1}\mathrm{d}g\in\mathfrak{g}$, valued in the Lie algebra $\mathfrak{g}$ of $G$. The flatness condition $\mathrm{d}j+\frac{1}{2}[j,j]=0$ is automatically satisfied. The equations of motion take the form
@@
where $v$ is a constant world-volume vector that explicitly breaks Lorentz invariance in $(1+2)$ dimensions.
Originally introduced by Manakov and Zakharov [7] for the case of $G=\mathrm{SU}(2)$, the model was shown to be solvable by the inverse scattering transform. It supports soliton solutions localised in all spatial directions and does not show classical scattering, an identifying feature of integrability [8]. Ward [8] further demonstrated that for the model to have a well-defined energy functional, the vector $v$ must be space-like, although its integrability remains unaffected by the nature of $v$. The MZW-model possesses an infinite hierarchy of conserved charges arising from its Lax pair and zero-curvature formulation [8]. Although only the energy-momentum tensor yields local, finite Noether charges, Ioannidou and Ward constructed an infinite sequence of non-local but finite conserved charges using integrals of the currents and its powers [23]. These charges reflect the underlying loop-algebra and twistor-theoretic structure of the model [24], and parallel the hierarchy of integrable flows found in systems such as the matrix KP-hierarchy [25]. This integrable structure also governs soliton interactions and scattering phenomena [26].
Recently, the MZW-model has reemerged in the context of higher-dimensional Chern–Simons theory [27, 28], a unifying framework for integrable $\sigma$-models associated with Lie group symmetries. In this setting, the MZW-model appears as a special case in certain higher-dimensional generalisations [29, 30]. Partly because of its breaking of Lorentz invariance, it was not obvious in which physical settings the MZW-model appears. Moreover, it seemed unlikely that the MZW-model will play a role in membrane dynamics since it assumes a flat metric on $\Sigma$ in contrast to the dynamical induced metric of the bosonic membrane. However, this article will show that the MZW-model can be obtained as a post-Newtonian or large-wrapping decoupling limit of the bosonic membrane.
@@ 2 A decoupling limit for the membrane on T2
For this, let us consider the bosonic membrane in a toy model [^2] In the sense, that for now we do not focus on whether this is a solution to eleven-dimensional supergravity. background
$\mathcal{M}_{\tau}\times\mathcal{M}_{H}=(\mathbb{R}_{t}\times T^{2})\times\mathcal{M}_{H}$
supported by a background 4-form flux with on leg in the flat ’longitudinal’ space $\mathcal{M}_{\tau}$ and three in the ’transverse’ space $\mathcal{M}_{H}$.
Here, we present two (ultimately equivalent) views on a decoupling limit that can be interpreted either as a large wrapping limit around the $T^{2}$ or a non-relativistic expansion (presented in appendix A). In particular, we aim at the following, at leading order in these limits:
@@
  • membrane dynamics along $\mathcal{M}_{H}$ $\rightarrow$ MZW-model
@@
This will involve scaling limits – both on the target space and one on the membrane world-volume.
In the following, we focus on the decoupling of the internal gravity and that, at leading order, the 4-form flux decouples from the $\mathcal{M}_{\tau}$ membrane dynamics. In section 3 it is shown which concrete choice of $\mathcal{M}_{H}$ and $F$ leads to the MZW-model dynamics (1.4) and how this can be embedded into a supergravity solution.
For this, let consider a general ansatz introducing scales $R$ and $\bar{R}$ for the longitudinal ($\mathcal{M}_{\tau}$, coordinates [^3] In a slight abuse of notation, we take the same indices $K,L,M,$... for the transversal coordinates $\hat{X}^{M}$ and the world-volume coordinates $\sigma^{M}$. $\hat{X}^{M}$) and transversal spaces ($\mathcal{M}_{H}$, coordinates $\hat{X}^{m}$)
$\mathrm{d}s^{2}$ $=R^{2}\eta_{MN}\mathrm{d}\hat{X}^{M}\mathrm{d}\hat{X}^{N}+\bar{R}^{2}H_{mn}\mathrm{d}\hat{X}^{m}\mathrm{d}\hat{X}^{n},$ (2.1)
$F$ $=\frac{1}{3!}R\bar{R}^{2}F_{klmN}\mathrm{d}\hat{X}^{k}\wedge\mathrm{d}\hat{X}^{l}\wedge\mathrm{d}\hat{X}^{m}\wedge\mathrm{d}\hat{X}^{N}.$ (2.2)
The scaling of $F$ has to be introduced in order to be consistent with the supergravity equations – compare (3.3).
Moreover, on the world-volume we consider arbitrary fluctuations around a static wrapped membrane in a large wrapping number limit $\omega\rightarrow\infty$:
$=\left(\omega\sigma^{M}+X^{M}(\sigma),X^{m}(\sigma)\right).$
The parameter $\omega$ describes both a ’world-volume speed of light’ and the wrapping number around the torus. [^4] In principle, these parameters could be different, of course – i.e. $X^{M}=(\omega_{0}\sigma^{0},\omega_{1}\sigma^{1},\omega_{2}\sigma^{2})$. For the limit in question, we want to assume $\omega_{0}\sim\omega_{1}\sim\omega_{2}\sim\omega\rightarrow\infty$. Finite quotients could be compensated for by rescaling of the membrane coordinates $\sigma^{M}$. In the limit $\omega\rightarrow\infty$, the $\mathcal{O}(\omega^{0})$ world-volume fields $X^{\mu}(\sigma)$ are suppressed and the induced metric simplifies significantly. In particular, one can compute
$\gamma^{MN}$ $=\frac{1}{\omega^{2}R^{2}}\left(\eta^{MN}-\frac{1}{\omega}2\partial^{(M}X^{N)}+\mathcal{O}(\omega^{-2})\right),$ (2.4)
$\sqrt{-\gamma}$ $=\omega^{3}R^{3}\left(1+\frac{1}{\omega}\partial_{M}X^{M}+\mathcal{O}(\omega^{-2})\right),$
in order to compute the expansion of the equations of motion (1.3) for $X^{M}$ and $X^{M}$ at leading order
$0$ $=\omega R\left(0+\mathcal{O}\left(\omega^{-1}\right)\right)$ (2.5)
${}\quad+\frac{\bar{R}^{2}}{R}\frac{1}{3!}\eta^{KM}F_{Mklm}\partial_{N}X^{k}\partial_{L}X^{l}\partial_{R}X^{m}\epsilon^{NLR},$
[…]
@@
${}\quad\left.+\frac{1}{2}H^{mn}F_{klnN}\partial_{K}X^{k}\partial_{L}X^{l}\epsilon^{KLN}+\mathcal{O}\left(\omega^{-1}\right)\right).$
Besides $\omega\rightarrow\infty$, the following scaling for the (dimensionless) scales $R$ and $\bar{R}$ is required, simultaneously ensuring the form (2.4) and guaranteeing that the transversal fluctations decouple from from the four-form flux at leading order:
$\frac{\bar{R}^{2}}{R^{2}\omega}\rightarrow 0.$ (2.7)
Remarkably, at leading order in the large wrapping limit the contributions from the metric fluctations (2.4) cancel completely in the equations motion (2.5).
There are several ways to interpret this limit. Let us outline two interpretations with very different consequences:
@@
  • One can solve (2.7) with the non-Lorentzian parameterisation introduced in appendix A. For this let us introduce the dimensionless scale $c$ with
    $R^{2}=c^{2},\qquad\bar{R}^{2}=\frac{1}{c}.$ (2.8)
@@
    In order to satisfy (2.7), the scaling of $\omega\sim c^{\alpha}\rightarrow\infty$ has to be chosen accordingly in terms of $c$. There are two choices
@@
    • a hyper-relativistic (Carollian) limit $c\rightarrow 0$: $\alpha<-3$.
    Interestingly, in both cases this corresponds to large energy limits $E=\omega R\rightarrow\infty$. Although at leading order one reproduces the MZW-model, these cases will lead to additional equations of motion at the finite but subleading order. [^5] Additional contributions would arise at order $\mathcal{O}(R)<\mathcal{O}(\omega R)$ or $\mathcal{O}(\bar{R}^{2}/R)\mathcal{O}(\omega R)$.
  • One can require that all the subleading contributions are supressed, and only the leading order contributions (proportional to $\omega R$) could be potentially finite. This leads to the following additional set of constraints.
    $R\rightarrow 0,\quad\frac{\bar{R}^{2}}{R}\rightarrow 0$ (2.9)
    This is solved by a parameterisation $R\sim\bar{R}\sim\frac{1}{\omega^{\alpha}}$, with $\alpha>0$. Then this limit can be interpreted as either a large or small energy $\omega R$ limit for $\alpha>0$ or $\alpha<0$ respectively. In any case, this is a small volume regime in target space – both scales $R,\bar{R}\rightarrow 0$.

### v3

@@ 1.1 The bosonic membrane
The action of the bosonic membrane (or equivalently the bosonic part of the supermembrane [2]) in a bosonic background of 11d supergravity, described by a metric $g$ and three-form gauge field $C$ is
$S$ $=T\int\mathrm{d}^{3}\sigma\left(\sqrt{-\gamma}+C_{3}\right)$ (1.1)
in its Nambu-Goto type form with $\gamma=\text{det}(\gamma_{MN})$ and
$\gamma_{MN}$ $=\partial_{M}\hat{X}^{\mu}\partial_{N}\hat{X}^{\nu}g_{\mu\nu},$
$C_{3}$ $=\frac{1}{3!}\epsilon^{KLM}\partial_{K}\hat{X}^{\mu}\partial_{L}\hat{X}^{\nu}\partial_{M}\hat{X}^{\rho}C_{\mu\nu\rho}.$
$\sigma^{M}$ are the coordinates on the membrane world-volume, $\hat{X}^{\mu}(\sigma)$ the embedding coordinate fields, and $T$ the membrane tension.
Due to the square-root, the equations of motion are highly non-linear and quite intractable. Even a Polyakov-like formulation, presented in appendix B, is not advantageous here as the world-volume metric cannot be gauge-fixed to conformally flat. Famously [17, 18], even this supposedly most simple case leads to complications, such as a continuous spectrum and spike instabilities. Quantum non-integrability of the membrane in the sense of scattering of fluctuations around a classical solution, and its relation to string theory, has recently been shown in [19]. In this letter, we will see that a large wrapping limit together with a scaling limit of the background can simplify these equations. In the special case considered here, the result will even be a classically integrable system.
@@ 1.2 The Manakov-Zakharov-Ward model
The model is of $\sigma$-model type and can be formulated in terms of maps $g:\Sigma\rightarrow G$ from a $(1+2)$-dimensional flat spacetime $\Sigma$ with standard coordinates $\sigma^{M}$ (referred to as the world-volume) into a Lie group $G$. The dynamics is conveniently expressed using the left-invariant Maurer–Cartan current $j=g^{-1}\mathrm{d}g\in\mathfrak{g}$, valued in the Lie algebra $\mathfrak{g}$ of $G$. The flatness condition $\mathrm{d}j+\frac{1}{2}[j,j]=0$ is automatically satisfied. The equations of motion take the form
@@
where $v=\mathrm{d}\sigma^{y}$ – with $\sigma^{y}$ denoting a distinguished coordinate for $y=0,1$ or $2$ – is a constant world-volume one-form with unit norm. The appearance of $v$ in the equations of motion manifestly breaks Lorentz invariance of this model in $(1+2)$ dimensions. Let us also present an action that gives rise to the equations of motion (1.2)
$S_{MZW}=T_{0}\int\left(\frac{1}{2}\text{tr}(j\wedge\star j)+\frac{1}{3}\sigma^{y}\ \text{tr}(j\wedge[j\overset{\wedge}{,}j])\right)$ (1.3)
where tr denotes some ad-invariant non-generate bilinear form on $\mathfrak{g}$. To the knowledge of the author, this action has not been presented in the literature before.
Originally introduced by Manakov and Zakharov [7] for the case of $G=\mathrm{SU}(2)$ and a $v$ being timelike, $v=\mathrm{d}\sigma^{0}$, the model was shown to be solvable by the inverse scattering transform. Ward [8] demonstrated that for the model to have a well-defined energy functional, the vector $v$ must be space-like, although its integrability remains unaffected by the nature of $v$. Depending on the nature of $v$ the model remains with $(1+1)$-dimensional Lorentz or two-dimensional Euclidean symmetry. Choosing $v=\mathrm{d}\sigma^{2}$ and coordinates222with light-cone coordinates $\sigma^{\pm}=\frac{1}{2}(\sigma^{0}\pm\sigma^{1})$ $(\sigma^{+},\sigma^{-},\sigma^{y}$), the equations of motion can be written explicitly as
$\partial_{-}j_{+}=\partial_{y}j^{y}.$ (1.4)
These equations can be obtained as the integrability condition of the auxiliary linear system [8]
$A(\sigma)\tilde{g}(\sigma,z)$ $=(z\partial_{y}-\partial_{+})\tilde{g}(\sigma,z)$ (1.5)
$B(\sigma)\tilde{g}(\sigma,z)$ $=(z\partial_{-}-\partial_{y})\tilde{g}(\sigma,z)$ (1.6)
with a complex spectral parameter $z$ and appropriate reality conditions. The group valued field $g$ is reproduced as $g(\sigma)=\tilde{g}(\sigma,0)$. Using this overdetermined linear system, one can construct soliton solutions. The MZW model supports soliton solutions localised in all spatial directions and does not show classical scattering for lump solitons, an identifying feature of integrability [8]. Ioannidou and Ward constructed an infinite sequence of non-local but finite conserved charges using integrals of the currents and its powers [23]. These charges reflect the underlying loop-algebra and twistor-theoretic structure of the model [24], and parallel the hierarchy of integrable flows found in systems such as the matrix KP-hierarchy [25]. This integrable structure also governs soliton interactions and scattering phenomena [26].
Recently, the MZW-model has reemerged in the context of higher-dimensional Chern–Simons theory [27, 28], a unifying framework for integrable $\sigma$-models associated with Lie group symmetries. In this setting, the MZW-model appears as a special case in certain higher-dimensional generalisations [29, 30]. Partly because of its breaking of Lorentz invariance, it was not obvious in which physical settings the MZW-model appears. Moreover, it seemed unlikely that the MZW-model will play a role in membrane dynamics since it assumes a flat metric on $\Sigma$ in contrast to the dynamical induced metric of the bosonic membrane. However, this article will show that the MZW-model can be obtained as a large-wrapping decoupling limit of the bosonic membrane.
@@ 2 A decoupling limit for the membrane on T2
For this, let us consider the bosonic membrane in a toy model background – in the sense, that for now we do not focus on whether this is a solution to eleven-dimensional supergravity:
$\mathcal{M}_{\tau}\times\mathcal{M}_{H}=(\mathbb{R}_{t}\times T^{2})\times\mathcal{M}_{H}$ (2.1)
supported by a certain background 4-form flux. We make the following ansatz for the metric and three-form gauge field, introducing scales $R$ and $\bar{R}$ for the longitudinal ($\mathcal{M}_{\tau}$, coordinates [^3] In a slight abuse of notation, we use the same indices $K,L,M,$... for the transversal coordinates $\hat{X}^{M}$ and the world-volume coordinates $\sigma^{M}$. $\hat{X}^{M}$) and transversal spaces ($\mathcal{M}_{H}$, coordinates $\hat{X}^{m}$)
$\mathrm{d}s^{2}$ $=R^{2}\eta_{MN}\mathrm{d}\hat{X}^{M}\mathrm{d}\hat{X}^{N}+\bar{R}^{2}H_{mn}(\hat{X}^{k})\ \mathrm{d}\hat{X}^{m}\mathrm{d}\hat{X}^{n},$ (2.2)
$F$ $=\frac{1}{3!}R\bar{R}^{2}F_{klmN}(\hat{X}^{n})\mathrm{d}\hat{X}^{k}\wedge\mathrm{d}\hat{X}^{l}\wedge\mathrm{d}\hat{X}^{m}\wedge\mathrm{d}\hat{X}^{N}.$
for a flat metric $\eta$ and a – for the purposes of this section – unspecified metric $H$ on $\mathcal{M}_{H}$. The scaling of $F$ has to be introduced in order to be consistent with the supergravity equations – compare (3.4) later. Moreover, we note the particular form of $F$ with one leg in $\mathcal{M}_{\tau}$ and the other in $\mathcal{M}_{H}$. In this section, we consider (2.2) as a toy model without considering it as a solution to the supergravity equations of motion.
On the world-volume, we assume arbitrary (finite) fluctuations around a static wrapped membrane in a large wrapping number limit $\omega\rightarrow\infty$:
$=\left(\omega\sigma^{M}+X^{M}(\sigma),X^{m}(\sigma)\right).$
The parameter $\omega$ describes both a ’world-volume speed of light’ and the wrapping number around the torus. [^4] In principle, these parameters could be different, of course – i.e. $X^{M}=(\omega_{0}\sigma^{0},\omega_{1}\sigma^{1},\omega_{2}\sigma^{2})$. For the limit in question, we want to assume $\omega_{0}\sim\omega_{1}\sim\omega_{2}\sim\omega\rightarrow\infty$. Finite quotients could be compensated for by rescaling of the membrane coordinates $\sigma^{M}$. In particular, we aim at the following, at leading order in this limit:
@@
This will involve scaling limits – both on the target space and on the membrane world-volume. In section 3 it is shown which concrete choice of $\mathcal{M}_{H}$ and $F$ leads to the MZW-model dynamics (1.2) and how this can be embedded into a supergravity solution.
The membrane action in the large wrapping limit.
In the limit $\omega\rightarrow\infty$, the $\mathcal{O}(\omega^{0})$ world-volume fields $X^{\mu}(\sigma)$ are suppressed, and the action (1.1) simplifies significantly. In particular,
$\gamma_{MN}$ $=\omega^{2}R^{2}\left(\eta_{MN}+\frac{1}{\omega}2\partial_{(M}X_{N)}\right$. (2.4)
${}\quad\left.+\frac{1}{\omega^{2}}\partial_{M}X^{K}\partial_{N}X^{L}\eta_{KL}+\frac{\bar{R}^{2}}{R^{2}\omega^{2}}\partial_{M}X^{k}\partial_{N}X^{l}H_{kl}\right),$
$\sqrt{-\gamma}$ $=\omega^{3}R^{3}\left(1+\frac{1}{\omega}\partial_{M}X^{M}+\right$. (2.5)
${}\quad+\frac{1}{\omega^{2}}\left(\partial_{K}X_{L}\partial^{K}X^{L}+(\partial_{K}X^{K})^{2}\right)$
${}\quad+\left.\frac{\bar{R}^{2}}{R^{2}\omega^{2}}\frac{1}{2}\eta^{MN}\partial_{M}X^{k}\partial_{N}X^{l}H_{kl}+\mathcal{O}\left(\frac{1}{\omega^{3}},\frac{\bar{R}^{4}}{R^{4}\omega^{4}}\right)\right).$
For this expansion to make sense, we also have to assume the following scaling:
$\frac{\bar{R}^{2}}{R^{2}\omega^{2}}\rightarrow 0.$ (2.6)
For writing the membrane action in this background, it is advantageous that $F$ in (2.2) is exact. The pull-back to the world-volume of its three-form potential $C$ can be written as
$C_{3}$ $=\omega R\bar{R}^{2}C^{H}_{3}+\mathcal{(}\omega^{0})$ (2.7)
with $C^{H}_{3}=-\frac{1}{3!}\omega R\bar{R}^{2}F_{klmN}(X^{n})\ \sigma^{N}\mathrm{d}X^{k}\wedge\mathrm{d}X^{l}\wedge\mathrm{d}X^{m}$. Note that the leading order does not depend on longitudinal fluctuations $X^{M}(\sigma)$.
The membrane action (1.1) at leading orders becomes
$S=\left(S_{H}[X^{k}]+\frac{R^{2}}{\bar{R}^{2}}S_{\tau}[X^{K}]+\mathcal{O}\left(\frac{1}{\omega},\frac{\bar{R^{2}}}{R^{2}\omega^{2}}\right)\right)$ (2.8)
in the limits $\omega\rightarrow\infty$ and $\frac{\bar{R}^{2}}{R^{2}\omega^{2}}\rightarrow 0$, with
$S_{H}$ $=T_{0}\int\mathrm{d}^{3}\sigma\left(\frac{1}{2}H_{mn}(X^{k})\partial_{M}X^{m}\partial_{N}X^{n}\ \eta^{MN}\right$.
@@
${}\qquad\qquad\left.-\frac{1}{3!}F_{klmN}(X^{n})\ \sigma^{N}\partial_{K}X^{k}\partial_{L}X^{l}\partial_{M}X^{m}\epsilon^{klm}\right),$
$S_{\tau}$ $=T_{0}\int\mathrm{d}^{3}\sigma\left(\partial_{K}X_{L}\partial^{K}X^{L}+(\partial_{K}X^{K})^{2}\right).$
The first two terms in (2.5) account for total derivatives in the action and will be neglected because they do not contribute to the equations of motion. In order to obtain a finite action, the following scaling of the membrane tension in terms of a finite $T_{0}$ was introduced:
$T\sim\frac{T_{0}}{\bar{R}^{2}R\omega}.$ (2.9)
$S_{H}$ depends only on the transverse and $S_{\tau}$ only on longitudinal fluctuations. Hence, as striven for, at leading orders we remain with models with trivial world-volume gravity (the world-volume metric becomes $\eta_{MN}$) and decoupling of transversal and longitudinal dynamics.
The fate of reparameterisation invariance.
At the level of the original membrane action, the model possesses reparameterisation invariance. In particular, one might be tempted to choose static gauge, $X^{M}\equiv 0$, for the longitudinal directions. In general, the theory after the limit (2.8), does not show reparameterisation invariance. Moreover, $S_{\tau}$ depends explicitly on longitudinal fluctations. It is not surprising that the resulting theories after the singular ’flat’ limit $\gamma\rightarrow\eta$ (at leading order) do not exhibit reparameterisation invariance and depend on the choice of parameterisation from where we start the limit.
Nevertheless, if one takes the scaling of $R$ and $\bar{R}$ as
$\frac{R}{\bar{R}}\rightarrow 0,$ (2.10)
interestingly, the leading order dynamics in (2.8) does not depend on the longitudinal fluctations, and hence this case is compatible with the choice of static gauge for the original membrane.
Interpretation of the limit.
There are several ways to choose interesting concrete realisations and interpret the combined scaling limits:
$\omega\rightarrow\infty,\quad\frac{\bar{R}}{R\omega}\rightarrow 0,\quad T\sim\frac{T_{0}}{\bar{R}^{2}R\omega}.$ (2.11)
@@
Let us outline one interpretation here. One can solve (2.11) with the non-Lorentzian parameterisation of the target space introduced in appendix A. For this, let us introduce the dimensionless scale $c$ with
$\omega=c^{\alpha},\quad R^{2}=c^{2},\quad\bar{R}^{2}=\frac{1}{c}.$ (2.12)
@@
In order to satisfy (2.11), the scaling of $\omega\sim c^{\alpha}\rightarrow\infty$ can be chosen accordingly in terms of $c$. There are two choices
@@
  • a hyper-relativistic (Carollian) limit $c\rightarrow 0$:
    $\alpha<-3/2$. Remarkably, this case leads to $\frac{R}{\bar{R}}\rightarrow 0$ and, therefore, to the total absence of the longitudinal fluctuations in (2.8).
Interestingly, in both cases this corresponds to a large energy ($E\sim\omega R\rightarrow\infty$) and small tension ($T\sim\frac{T_{0}}{\omega}\rightarrow 0$) limit. Nevertheless, many other ways of interpreting (2.11), with other physical consequences, are possible.